OpenStudy (anonymous):

The formula for degrees Celsius is: C = 5 over 9 (F − 32), where F stands for degrees Fahrenheit. Part 1: Solve the equation for F and show all steps. Part 2: Determine how many degrees Fahrenheit 20 degrees Celsius is. please help?

OpenStudy (blurbendy):

C = 5/9 (F – 32) 9C = 5(F - 32) F - 32 = 9C/5 F = 9C/5 + 32

OpenStudy (blurbendy):

Then put 20 in for C in the bottom equation

OpenStudy (anonymous):

it would be 4.86?